OpenStudy (anonymous):

For y = x2 + 6x - 16, Determine if the parabola opens up or down. State if the vertex will be a maximum or minimum. Find the vertex. Find the x-intercepts. Describe the graph of the equation.

OpenStudy (anonymous):

up

OpenStudy (anonymous):

as coeffi. of x^2 is positive , it opens up. hence the vertex will be minimum. vertex(-b/2a , -D/4a) x-int. ,put y=0.

OpenStudy (anonymous):

The Parabola opens up because the term with X^2 is positive If it was negative then it would open down.

OpenStudy (anonymous):

Yep, you're absolutely correct! Try plotting a few equations on here: http://www.mathsisfun.com/data/function-grapher.php You should get the hang of it in no time.

OpenStudy (kenljw):

y = x2 + 6x - 16 y = (x + 3)^2 - 9 - 16 y = (x + 3)^2 - 25 notice for all values of x > 2 y is positive therefore it must open up